Circular No. 7986

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002gd IN NGC 7537 Independent CCD discoveries of an apparent supernova in NGC 7537

have been received from A. Klotz, Guitalens, France, with a 0.20-m reflector (limiting mag ∼ 18.0) and from T. Puckett and A. Langoussis, Mountain Town, GA, with the Puckett Observatory 0.60-m automated supernova patrol telescope. Klotz provided the following approximate position for the new object: α = 23^{h}14^{m}36^{s}.5, δ = +4^{o}30′08′′ (equinox 2000.0), which is 34′′ east and 8′′ north of the nucleus of NGC 7537. Puckett provides the following precise-position end figures: 36^{s}.98, 05′′.7, which is 36′′.8 east and 10′′.8 north of the center of NGC 7537. W. M. Wood-Vasey, G. Aldering, and P. Nugent, Lawrence Berkeley National Laboratory, report the following precise-position end figures from a NEAT image taken at Palomar: 37^{s}.01, 06′′.1. Unfiltered CCD magnitude estimates (P = Puckett; W = Wood-Vasey; K = Klotz): 1951 Aug. 10 UT, [20.0: (Palomar Sky Survey, via Puckett); 2001 Sept. 6, [19.0 (W; S/N = 7); 2002 July 16, [20.5: (P); Aug. 13, [20.5: (P); Aug. 22, [20.5 (W; S/N = 7); Oct. 5.31, 18.9 (W); 5.33, 19.1 (W); 5.92, ∼ 16.7 (K); 6.22, 17.5 (P); 6.28, 18.3 (W); 6.31, 18.0 (W); 6.33, 18.1 (W); 6.83 ∼ 16.7 (K). Wood-Vasey notes that their preliminary magnitudes were obtained after subtraction of the galaxy light. Klotz adds that SN 2002gd is not present on numerous images taken by himself before during from 2000 Aug. 23 to 2002 Aug. 17. Puckett adds that SN 2002gd was also confirmed on CCD frames taken by R. Stephens, Rancho Cucamonga, CA (0.35-m reflector).

COMETS C/2002 S5, C/2002 S6, AND C/2002 S7 (SOHO) Further to IAUC 7984, D. Hammer reports two more comets found on

SOHO website C3 images by M. Oates (C/2002 S6, a Kreutz sungrazer) and R. Matson (C/2002 S7, a member of the Kracht group, also visible in the C2 coronagraph). The reduced astrometry and orbital elements by B. G. Marsden appear on MPEC 2002-T25, where R. Kracht was incorrectly stated as having found C/2002 S7.

Comet  2002 UT  α_{2000}  δ_{2000}
C/2002 S6 Sept.20.488  11^{h}32.^{m}9  + 0^{o}35′
C/2002 S7  20.721  11 39.9  + 2 21

Corrigendum. On IAUC 7984, last line (C/2002 S5), for [Sept.] 17.646 read 19.646

2002 October 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green